How much work is done on a bookshelf being pulled with a force of 53.8 N for a distance of 5.00 m at an angle of 37.0o from the horizontal?


1
Expert's answer
2022-01-21T13:48:41-0500

Given:

"F=53.8\:\rm N"

"d=5.00\:\rm m"

"\theta=37.0^\circ"


The work done

"W=Fd\cos\theta=\rm 53.8*5.00*\cos37^\circ=215\: J"